# LiDDA: Data Driven Attribution at LinkedIn

### v1

@@ Abstract.
Data Driven Attribution, which assigns conversion credits to marketing interactions based on causal patterns learned from data, is the foundation of modern marketing intelligence and vital to any marketing businesses and advertising platform. In this paper, we introduce a unified transformer-based attribution approach that can handle member-level data, aggregate-level data, and integration of external macro factors. We detail the large scale implementation of the approach at LinkedIn, showcasing significant impact. We also share learning and insights that are broadly applicable to the marketing and ad tech fields.
@@ 1. Introduction
Marketers rely on attribution measurement to understand the impact of marketing campaigns on customers’ journeys and decisions. Accurate attribution enables marketers to better optimize marketing spend and strategy to maximize their return on investment (ROI). In this paper, we propose a general framework for large scale end-to-end attribution system and details the implementation at LinkedIn, LiDDA. LiDDA leverages a combination of bottom-up and top-down modeling approaches in order to build a more accurate attribution platform to measure campaign performance and ROI. The system was originally developed for our internal Go-To-Market (GTM) marketing and has since expanded to support external applications for LinkedIn customers.
@@ 1.1. LinkedIn Use Case
GTM Marketing is responsible for managing digital marketing campaigns for various LinkedIn products (e.g., Recruiter Lite, Sales Navigator), targeting both acquisition and retention objectives, where audiences of marketing includes consumers and buyers. Consumer marketing focuses on the individual as a member such as member account signups. Buyer marketing focuses on the company-level conversions such as the use of LinkedIn Ads platform ads or Talent Solutions job posts.
@@
One major challenge is lack of a unified view of the member journey across multiple channels. Existing media-mix-modeling (MMM) approaches provided channel-level visibility; however they could not provide more detailed campaign breakouts and journey information. Last touch attribution (LTA) was the default standard; yet it biased results to bottom of funnel (BOFU) campaigns and omitted channels without individual identifiers.
@@ 1.2.2. Rules Based Attribution
Traditional attribution models uses rule-based attribution (RBA) approaches such as first-touch, last-touch, or time decay(Gaur and Bharti, 2020). These approaches assign conversion credit based on simple, predetermined rules at user level. RBA methods are easy to understand and implement however, they are often based on limited assumptions and introduces biases.
@@ 1.4. Our contributions
As far as we know, LiDDA is the first industry deployed large scale DDA system that leverage transformer-type attention mechanism and can handle member-level, aggregate-level and integration of external macro factors. In particular, the system is highlighted by: (i) Transformer Architecture: our use of transformer type of attention offered both single-pass inference and an explanability mechanis; (ii) Temporal-aware attention: we discretize time gaps between campaign touchpoints and learn positional embeddings—plus a special day‑of‑week embedding—to capture both sequence order and seasonal effects; (iii) Privacy‑driven data imputation: to correct for missing paid‑media signals (e.g. due to GDPR/CCPA restrictions), we probabilistically inject aggregated impressions back into user paths; (iv) Unified DDA&MMM outputs: we reconcile discrepancies between independent DDA and macro‑level MMM by scaling and aligning their channel attributions. (v) End‑to‑end solution: we present our modeling innovations, implementation details, comprehensive evaluation and insights from launching to both GTM marketing and Ads platform in the paper.
@@ 2.1. System Overview
LinkedIn Data Driven Attribution (LiDDA) is our modeled attribution platform (Figure  1) which utilizes an attention-based modeling approach. Attention models are well suited for this task since they retain sequential information about the buyer journey and also enable us to incorporate additional features about the member, their company, and campaign information.
Overall, the model is structured as a binary classification task to predict whether a buyer journey will result in a conversion
$P(C|E_{M},E_{C},S),$
where $C$ denote the binary conversion outcome, $E_{M}$ denote the member embedding with length $D_{M}$, $E_{C}$ denote the (member’s) company embedding with length $D_{C}$, and $S$ denote the sequence of marketing touchpoints on the member with length $N$. $E_{M}$ is a member embedding derived from LinkedIn data such as skills, titles, etc. that generally capture latent features of that member. Likewise, $E_{C}$ is a company embedding derived from LinkedIn data such as industry, roles, etc.
From the trained model, we output the aggregated attention weight matrix from the touchpoint sequence $S$. The attention weights are normalized such that they add up to 1.0, i.e.,
(1) $\hat{\alpha}_{i}=\frac{\alpha_{i}}{\sum_{j=1}^{n}\alpha_{j}}$
and the normalized weights are interpreted as percent contribution of that touchpoint to the final conversion event. We further discuss attention in Section 2.4.
@@ 2.2.1. Sequence Representation
The touchpoint sequence $S$ is a time series of touchpoints $T\in\{t_{0},\cdots,t_{K+1}\}$ indexed by timestep t. $K$ is the distinct number of possible (channel, action) pairs with 0 reserved for padding. An example of a (channel, action) could be $(EMAIL,OPEN)$ or $(AD,CLICK)$. The sequence $S$ is limited to the last $N$ touches such that paths with a length greater than $N$ have their earliest engagements truncated. The choice of touchpoints included are context-specific and could represent marketing activities or other attributable steps in a defined journey.
@@ 2.2.2. Touchpoint Representation
An additional feature, $E_{MCID}$, is concatenated to the learned $E_{T}$ at each timestep. $E_{MCID}\in\mathbb{R}^{MCID\times{\text{dim}}}$ is an embedding representation of our marketing campaign which captures different aspects of the campaign such as its metadata or content. We obtain these embeddings from an upstream model which extracts a text description of the campaign and generates an embedded representation using an internal LLM fine-tuned on a LinkedIn corpus. This approach allows flexibility in combining structured and unstructured data consistently across domains.
@@ 2.3.1. Positional Encodings
Attention models do not inherently model the order of touches in the paths so additional encodings are needed to capture the relative positions of each touch. Sinusoidal positional encodings are commonly used for temporal ordering as in the original attention design (Vaswani et al., 2017). These are implemented as:
(2) $PE_{(pos,2i)}=sin(pos/10000^{2i/d_{model}})$
$PE_{(pos,2i+1)}=cos(pos/10000^{2i/d_{model}})$
Because our touchpoint embedding $E_{T}$ is at the granularity of (channel, action), the number of possible combination is far lower than typical NLP vocabulary sizes. Therefore we use a substantially smaller embedding dimension for our data. We employ Time Absolute Position Encodings (tAPE) where the positional encoding definition in Equation (2) is modified to adjust for the lower dimensional representation(Foumani et al., 2023). The tAPE encodings are defined as:
(3) $\omega_{k}=10000^{-2k/d_{model}}$
$\omega_{k}^{new}=\frac{\omega_{k}\times d_{model}}{L}$
@@ 2.3.2. Positional Embeddings
We combine all the embeddings together for each touchpoint as:
(4) $concat(E_{MCID},E_{T})+E_{D}+{E_{DOW}}+tAPE$
@@ 2.4.1. Attention for Attribution
Prior works (li et al., 2018; Kumar et al., 2020) utilize attention-based crediting by applying a self-attention layer over an recurrent network. Our initial exploration of this approach found a heavy bias towards end-of-path touches. Using a bi-directional LSTM may help alleviate this effect, however we modified the approach to directly apply self attention over the input sequence. One issue with the LSTM encoder is that the hidden state contains information about previous timesteps, making attention interpretation less straight forward (Serrano and Smith, 2019). The topic of attention interpretability has been the subject of extensive research (Jain and Wallace, 2019; Wiegreffe and Pinter, 2019). The central concern is that the resulting weights may not faithfully align with the model final output. While our modification can still introduce interpretability issues in attention based models (Brunner et al., 2020), we observe within our domain that attention weights do impact model output. We share additional findings in Appendix A.1.
It is also possible to compute attribution based on the incremental propensity from the prior engagement (Shao and Li, 2011) or as an approximation of Shapely values (Yao et al., 2022). These methods are not necessarily superior. Counterfactual simulations can be computationally expensive and require accurately determining the counterfactual which is non-trivial. We did evaluate the attributions under the attention vs incremental propensity and found negligible differences (Table 1). We also observed that incremental propensity approach often produces early jumps causing front-loaded attribution as propensity gains saturate with length.
@@
Table 1. A LiDDA trained model was used to perform attribution under two methods. $Incremental$ assigns attribution according to the incremental change in $P(C|T_{i})-P(C|T_{i-1})$ while $Attention$ uses only the attention weights.
@@ 2.5. Imputation of Paid Media Touchpoints
LinkedIn’s outbound marketing channels are grouped into owned and paid channels. When marketers use owned channels, Email and LinkedIn’s advertising platform, to market products and services to members, we can collect touchpoints from all actions at the user level, e.g., sends, opens, clicks, impressions. On the other hand, when external (paid) channels are used, due to privacy regulations like GDPR or CCPA, we cannot readily collect data at the user level (Figure  3).

### v2

@@ Abstract.
Data Driven Attribution, which assigns conversion credits to marketing interactions based on causal patterns learned from data, is the foundation of modern marketing intelligence and vital to any marketing businesses and advertising platform. In this paper, we introduce a unified transformer-based attribution approach that can handle member-level data, aggregate-level data, and integration of external macro factors. We detail the large scale implementation of the approach at LinkedIn, showcasing significant impact. We also share learnings and insights which are broadly applicable to the marketing and ad tech fields.
@@ 1. Introduction
Marketers rely on attribution measurement to understand the impact of marketing campaigns on customers’ journeys and decisions. Accurate attribution enables marketers to better optimize marketing spend and strategy to maximize their return on investment (ROI). In this paper, we propose a framework for large scale end-to-end attribution system and details the implementation at LinkedIn. LinkedIn Data Driven Attribution (LiDDA) leverages a combination of bottom-up and top-down modeling approaches in order to build a more accurate attribution platform to measure campaign performance and ROI. The system was originally developed for our internal Go-To-Market (GTM) Marketing and has since expanded to support external applications for LinkedIn customers.
@@ 1.1. LinkedIn Use Case
GTM Marketing is responsible for managing digital marketing campaigns for various LinkedIn products (e.g., Recruiter Lite, Sales Navigator), targeting both acquisition and retention objectives, where marketing audiences include both consumers and buyers. Consumer marketing focuses on the individual as a member such as member account signups. Buyer marketing focuses on the company-level conversions such as the use of LinkedIn Ads platform ads or Talent Solutions job posts.
@@
One major challenge is lack of a unified view of the member journey across multiple channels. Existing media-mix-modeling (MMM) approaches provide channel-level visibility; however they lack detailed campaign breakouts and journey information. Last touch attribution (LTA) was the default standard; yet it biased results to bottom of funnel (BOFU) campaigns and omitted channels without individual identifiers.
@@ 1.2.2. Rules Based Attribution
Traditional attribution models commonly use rule-based attribution (RBA) approaches such as first-touch, last-touch, or time decay(Gaur and Bharti, 2020). These approaches assign conversion credit based on simple, predetermined rules at user level. RBA methods are easy to understand and implement however, they are often based on limited assumptions and introduces biases.
@@ 1.4. Our contributions
As far as we know, LiDDA is the first industry deployed large scale DDA system that leverage transformer-type attention mechanism and can handle member-level, aggregate-level and integration of external macro factors. In particular, the system is highlighted by: (i) Transformer Architecture: our use of transformer type of attention offered both single-pass inference and an explanability mechanis; (ii) Temporal-aware attention: we discretize time gaps between campaign touchpoints and learn positional embeddings—plus a special day‑of‑week embedding—to capture both sequence order and seasonal effects; (iii) Privacy‑driven data imputation: to correct for missing paid‑media signals (e.g. due to GDPR/CCPA restrictions), we probabilistically inject aggregated impressions back into user paths; (iv) Unified DDA&MMM outputs: we reconcile discrepancies between independent DDA and macro‑level MMM by training-time calibration; (v) End‑to‑end solution: we present our modeling innovations, implementation details, comprehensive evaluation and insights from launching to both GTM marketing and Ads platform in the paper.
@@ 2.1. System Overview
LiDDA is our modeled attribution platform (Figure  1) which utilizes an attention-based modeling approach. Attention models are well suited for this task since they retain sequential information about the buyer journey and also enable us to incorporate additional features about the member, their company, and campaign information.
Overall, the model is structured as a binary classification task to predict whether a buyer journey will result in a conversion $P(Y|E_{M},E_{C},S),$ where $Y$ denotes the binary conversion outcome, $E_{M}$ denotes the member embedding with length $D_{M}$, $E_{C}$ denotes the (member’s) company embedding with length $D_{C}$, and $S$ denotes the sequence of marketing touchpoints on the member with length $N$. $E_{M}$ is a member embedding derived from LinkedIn data such as skills, titles, etc. that generally capture latent features of that member. Likewise, $E_{C}$ is a company embedding derived from LinkedIn data such as industry, roles, etc.
From the trained model, we output the aggregated attention weight matrix from the touchpoint sequence $S$. The attention weights are normalized such that they add up to 1.0, i.e., $\hat{\alpha}_{i}=\frac{\alpha_{i}}{\sum_{j=1}^{n}\alpha_{j}},$ and the normalized weights are interpreted as percent contribution of that touchpoint to the final conversion event. We further discuss attention in Section 2.4.
@@ 2.2.1. Sequence Representation
The touchpoint sequence $S$ is a time series of touchpoints $t\in\mathcal{T}$. $K=|\mathcal{T}|$ is the distinct number of possible (channel, action) pairs with 0 reserved for padding. An example of a (channel, action) could be $(EMAIL,OPEN)$ or $(AD,CLICK)$. The sequence $S$ is limited to the last $N$ touches such that paths with a length greater than $N$ have their earliest engagements truncated. The choice of touchpoints included are context-specific and could represent marketing activities or other attributable steps in a defined journey.
@@ 2.2.2. Touchpoint Representation
An additional feature, $E_{MCID}$, is concatenated to the learned $E_{T}$ at each timestep. The embedding $E_{MCID}\in\mathbb{R}^{MCID\times{\text{dim}}}$ is a representation of our marketing campaign which captures different aspects of the campaign such as its metadata or content. We obtain these embeddings from an upstream model which extracts a text description of the campaign and generates an embedded representation using an internal LLM fine-tuned on a LinkedIn corpus. This approach allows flexibility in combining structured and unstructured data consistently across domains.
@@ 2.3.1. Positional Encodings
Attention models do not inherently model the order of touches in the paths so additional encodings are needed to capture the relative positions of each touch. Sinusoidal positional encodings are commonly used for temporal ordering as in the original attention design (Vaswani et al., 2017). These are implemented as: $PE_{(pos,2i)}=sin(pos/10000^{2i/d_{model}})$, $PE_{(pos,2i+1)}=cos(pos/10000^{2i/d_{model}})$
Because our touchpoint embedding $E_{T}$ is at the granularity of (channel, action), the number of possible combination is far lower than typical NLP vocabulary sizes. Therefore we use a substantially smaller embedding dimension for our data. We employ Time Absolute Position Encodings (tAPE) where the positional encoding above is modified to adjust for the lower dimensional representation(Foumani et al., 2023). The tAPE encodings are defined as: $\omega_{k}=10000^{-2k/d_{model}}$, $\omega_{k}^{new}=\frac{\omega_{k}\times d_{model}}{L}$.
@@ 2.3.2. Positional Embeddings
We combine all the embeddings together for each touchpoint as: $concat(E_{MCID},E_{T})+E_{D}+{E_{DOW}}+tAPE$.
@@ 2.4.1. Attention for Attribution
Prior works (li et al., 2018; Kumar et al., 2020) utilize attention-based crediting by applying a self-attention layer over an recurrent network. Our initial exploration of this approach found a heavy bias towards end-of-path touches. Using a bi-directional LSTM may help alleviate this effect, however we modified the approach to directly apply self attention over the input sequence. One issue with the LSTM encoder is that the hidden state contains information about previous timesteps, making attention interpretation less straight forward (Serrano and Smith, 2019). The topic of attention interpretability has been the subject of extensive research (Jain and Wallace, 2019; Wiegreffe and Pinter, 2019). The central concern is that the resulting weights may not faithfully align with the model final output. While our modification can still introduce interpretability issues in attention based models (Brunner et al., 2020), we observe within our domain that attention weights do impact model output. We share additional findings in Appendix B.2.
It is also possible to compute attribution based on the incremental propensity from the prior engagement (Shao and Li, 2011) or as an approximation of Shapely values (Yao et al., 2022). These methods are not necessarily superior. Counterfactual simulations can be computationally expensive and require accurately determining the counterfactual which is non-trivial. We did evaluate the attributions under the attention vs incremental propensity and found negligible differences (Table 1). We also observed that incremental propensity approach often produces early jumps causing front-loaded attribution as changes in propensity saturate with length.
@@
Table 1. Comparison of incremental propensity and attention weight from LiDDA model.
@@ 2.5. Imputation of Paid Media Touchpoints
LinkedIn’s outbound marketing channels are grouped into owned and paid channels. For owned channels, we can collect touchpoints from all actions at the user level, e.g., sends, opens, clicks, impressions, while for external channels, impression data at the user level are not available, e.g., due to privacy regulations like GDPR or CCPA (Appendix Figure 6).
Algorithm 1 Allocate Member Impressions
Inputs:
 External $\{I^{\rm ext}_{c,d},C^{\rm ext}_{c,d}\}$, clicks $J_{i,c,d}$
 Owned Impressions $\{I^{\rm own}_{\rm total},I^{\rm own}_{\rm clickers}\}$
 No-click paths $\mathcal{P}_{d}$, path‐weights $\{u_{k}\}$
Outputs: click‐imps $\{m_{j}\}$, path‐imps $\{I_{k}\}$
Step 1: Member‐only budget
$I_{d}=\sum_{c}I^{\rm ext}_{c,d}\;\frac{\sum_{i}|J_{i,c,d}|}{C^{\rm ext}_{c,d}}$
Step 2: Stochastic click allocation
$\alpha=\frac{I^{\rm own}_{\rm clickers}}{I^{\rm own}_{\rm total}},\quad r_{c,d}=\alpha\,\frac{I^{\rm ext}_{c,d}}{C^{\rm ext}_{c,d}}$
$m_{j}\;\sim\;\mathrm{Geom}\bigl{(}p=\tfrac{1}{r_{c(j),d}}\bigr{)},\quad L_{d}=I_{d}-\sum_{j}m_{j}$
Step 3: Stochastic leftover allocation
$p_{k}=\frac{u_{k}}{\sum_{\ell\in\mathcal{P}_{d}}u_{\ell}},\quad\lambda_{k}=L_{d}\,p_{k},\quad\tilde{I}_{k}\sim\mathrm{Poisson}(\lambda_{k})$
$I_{k}=\Bigl{\lfloor}\tilde{I}_{k}\,\frac{L_{d}}{\sum_{\ell}\tilde{I}_{\ell}}\Bigr{\rfloor}$